### 14 Mar 2026 v7

@@ 5.3.1 $Eagle^{*}$ key generator
The key is divided into three independent variables, denoted as $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$, where $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$ both have odd number of bits with 1.
@@ 5.3.2 $Eagle^{*}$ encryption algorithm
We randomly generate $2(K+3)$ groups number of $L$ bits, denoted as $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, where $x_{i}$ and $y_{i}$ have odd number of bits different.
Then we randomly generate another four number as $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
@@
0:Keys: $(\delta_{1},\delta_{2},\delta_{3})$. Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$. Random number sequences: $x=(x_{1},x_{2},...,x_{K+3})$,                $y=(y_{1},y_{2},...,y_{K+3})$.Random number: $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
0:Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3}$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$
1:for $i=1;i\leq K+3;i++$ do
2:  $Sx_{i}\leftarrow x_{i}*\delta_{1}*\delta_{1}$
3:  $Sy_{i}\leftarrow y_{i}*\delta_{2}*\delta_{2}$
4:end for
5:for $j=1;j\leq K+3;j++$ do
6:  $(C_{j},S_{j})\leftarrow\mathbb{E}_{x_{j},y_{j}}(P_{j},S_{j-1})$
7:  if $j<K+3$ then
8:   $S_{j}\leftarrow S_{j}*\delta_{3}*\delta_{3}$
9:  end if
10:end for
@@
$S_{0}$$x_{1}*\delta_{1}^{2}=Sx_{1};y_{1}*\delta_{2}^{2}=Sy_{1}$$S_{1}*\delta_{3}^{2}$$C_{1}$$x_{2}*\delta_{1}^{2}=Sx_{2};y_{2}*\delta_{2}^{2}=Sy_{2}$$S_{2}*\delta_{3}^{2}$$C_{2}$$S_{K+2}*\delta_{3}^{2}$$C_{K+2}$$x_{K+3}*\delta_{1}^{2}=Sx_{K+3};y_{K+3}*\delta_{2}^{2}=Sy_{K+3}$$S_{K+3}$$C_{K+3}$$\scriptstyle{P_{1},x_{1},y_{1}}$$\scriptstyle{P_{2},x_{2},y_{2}}\scriptstyle{......}$$\scriptstyle{P_{K+3},x_{K+3},y_{K+3}}$
@@ 5.3.3 $Eagle^{*}$ decryption algorithm
For $Eagle^{*}$ decryptors, the ciphertext sequences are $C=($$C_{1}$, $C_{2}$, …, $C_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ and $S_{K+3}$. Here, the ciphertext sequences can be divided into three independent parts $(C_{1},C_{2}$, …, $C_{K+3},S_{K+3})$ and $(Sx_{1},Sx_{2}$, …, $Sx_{K+3})$ and $(Sy_{1},Sy_{2}$, …, $Sy_{K+3})$, First we use $(\delta_{1})$ and $(Sx_{1},Sx_{2},...,Sx_{K+3})$ to calculate the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, then we use $(\delta_{2})$ and $(Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ to calculate the random number sequences $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, then we use the key $(\delta_{3})$ and the ciphertexts $(C_{1}$, $C_{2}$, …, $C_{K+3}$, $S_{K+3})$ and the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$ and $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$ to decrypt $P=(P_{1}$, $P_{2}$, …, $P_{K})$. The specific process is detailed in the following algorithm.
@@
0:Key: $(\delta_{1},\delta_{2},\delta_{3})$. Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3},$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$.
@@
1:for $i=1;i\leq K+3;i++$ do
2:  $x_{i}\leftarrow Sx_{i}/\delta_{1}/\delta_{1}$
3:  $y_{i}\leftarrow Sy_{i}/\delta_{2}/\delta_{2}$
4:end for
5:for $j=K+3;j\geq 1;j$-- do
6:  $(P_{j},S_{j-1})\leftarrow\mathbb{D}_{x_{j},y_{j}}(C_{j},S_{j})$
7:  $S_{j-1}\leftarrow S_{j-1}/\delta_{3}/\delta_{3}$
8:end for
@@ 5.3.4 Basic Security Analysis
[Theorem 11] In the $Eagle^{*}$ encryption algorithm, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to cracking two respective keys $\delta_{1}$ and $\delta_{2}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm and only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
@@
According to the definition of $Eagle^{*}$, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to solve the following system of equations.
@@
Where $x_{1},x_{2},...,x_{K+3},y_{1},y_{2},...,y_{K+3},P_{K+1},P_{K+2},P_{K+3},\delta_{1},\delta_{2}$ are unknown variables to be solved. $P_{1}$, $P_{2}$, …, $P_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$, $S_{K+3}$ are known variables.
@@
The unknown variables $x_{1}$, $x_{2}$, …, $x_{K+3}$ in the system of equations (5.6) can be equivalently regarded as the plaintexts of the A-encryption algorithm, with the known variables $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$ being the ciphertexts of the A-encryption algorithm, and $\delta_{1}$ being the key. Solving the system of equations (5.6) is equivalent to crack the key $\delta_{1}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm.
Similarly, solving the system of equations (5.7) is equivalent to crack the key $\delta_{2}$ under the condition of knowing only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
@@
(ii) In the case of knowing only ciphertext, any key for one group is equally possible.
(iii) In the case of knowing only ciphertext, any plaintext in one group is equally possible.
According to the conclusion of Theorem 7, A-Encrypt obviously satisfies the above three conditions. This means that A-Encrypt can only use exhaustive search to crack the key when only knowing their ciphertexts, and there are no effective algorithms.
@@
It suffices to consider only the case where $\delta_{1}$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6), but this does not change the conclusion that each equation in the system of equations (5.5) is an identity equation.
@@
When only exhaust $\delta_{3}$, since both $x_{i}$ and $y_{i}$ are unknown, it is impossible to determine $S_{i}$ solely based on $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
@@
When enumerating over $(\delta_{1},\delta_{2})$, we can directly calcute $x_{1}$, $x_{2}$, …, $x_{K+3}$ from the system of equations (5.6), and directly calculate $y_{1}$, $y_{2}$, …, $y_{K+3}$ from the system of equations (5.7). Substituting these calculated $x_{i}$ and $y_{i}$ into the system of equations (5.5), we can sequentially calculate the hidden variables $S_{1}$, $S_{2}$, …, $S_{K+2}$. Then we can solve the equation $S_{i}*\delta_{3}*\delta_{3}=S_{i+1}$ to obtain $\delta_{3}$, this equation is a nonlinear equation about $\delta_{3}$. As long as Assertion 1 holds, the only way to solve this equation is exhaustively enumerating $\delta_{3}$.
@@
It suffices to consider only the case where $(\delta_{1},\delta_{3})$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6). Since $y_{i}$ is unknown, it is impossible to determine $S_{i}$ solely based on $x_{i}$ and $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
@@
Due to the complete randomness of $x_{i}$ and $y_{i}$, there is no statistical relationship or any form of mathematical relationship between the plaintext $P_{i}$ and $Sx_{i}$. Similarly, there is no statistical relationship between $P_{i}$ and $Sy_{i}$. Furthermore, according to Theorem 3, $P_{i}$ and $C_{i}$ are also completely unrelated. In fact, if we disregard $S_{i}$ and $\delta_{3}$ (in reality, only $S_{K+3}$ and $\delta_{3}$ are known, all $S_{i}$ are invisible without solving for $x_{i}$ and $y_{i}$), the encryption process of $Eagle^{*}$ is equivalent to one-time pad. An attacker cannot find any patterns between plaintexts and ciphertexts from a large sample of plaintext-ciphertext pairs.
Lastly, it should be noted that during encryption, why do we need to randomly generate three additional sets of plaintexts $(P_{K+1},P_{K+2},P_{K+3})$ after $P_{K}$? This is because there exists an implicit relationship between $S_{K+3}$ and $P_{K+3}$ in the ciphertext, which is $S_{K+3}=P_{K+3}*(x_{K+3}\oplus y_{K+3})\oplus x_{K+3}^{+(L-1)}$. In order to make this implicit relationship invisible, we generated several additional sets of plaintext sequences when designing $Eagle^{*}$. The reason why we chose three sets is that there are three variables in the key, so we opted to generate three additional sets of plaintexts.

### 06 Apr 2026 v8

@@ 6.1.1 $Eagle^{*}$ key generator
The key is divided into three independent variables, denoted as $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$, where $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$ both contain odd number of bits that are 1.
@@ 6.1.2 $Eagle^{*}$ encryption algorithm
We randomly generate $3K+2$ groups number with each have $L$ bits, denoted as $x=(x_{1}$, $x_{2}$, …, $x_{K+1})$, $y=(y_{1}$, $y_{2}$, …, $y_{K+1})$, $z=(z_{1}$, $z_{2}$, …, $z_{K})$, where $x_{i}$ contains even number of bits that 1, $y_{i}$ contains odd number of bits that are 1, $z_{i}$ is odd number.
Then we randomly generate three number as $u$, $M_{0}$ and $P_{K+1}$, where $u$ is an odd number which contain odd number of bits with 1, $M_{0}$ is an odd number, $P_{K+1}$ is an odd number.
@@
0:Keys: $(\delta_{1},\delta_{2},\delta_{3})$. Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$. Random number sequences:                 $x=(x_{1},x_{2},...,x_{K+1})$,                $y=(y_{1},y_{2},...,y_{K+1})$.                $z=(z_{1},y_{2},...,z_{K})$.Random number: $u,M_{0},P_{K+1}$.
0:Ciphertext sequences:                 $(C_{1},C_{2},...,C_{K+1})$,                 $(Sx_{1},Sx_{2},...,Sx_{K+1})$,                 $(Sy_{1},Sy_{2},...,Sy_{K+1})$,                 $(Sz_{1},Sz_{2},...,Sz_{K})$,                 $Su,P_{K+1}$
[…]
2:for $i=1;i\leq K+1;i++$ do
3:  $Sx_{i}\leftarrow(x_{i}\oplus 1)*\delta_{1}*\delta_{1}$
4:  $Sy_{i}\leftarrow y_{i}*\delta_{2}*\delta_{2}$
5:end for
[…]
8:for $j=1;j\leq K;j++$ do
9:  $Sz_{j}\leftarrow P_{j}\oplus z_{j}$
10:  $(C_{j},C)\leftarrow(z_{j},M)\bigotimes(x_{j},y_{j})$
11:  $M\leftarrow C*u*u$
12:  $M\leftarrow\phi(M)$
13:end for
[…]
@@
$M_{0}$$(z_{1},M_{0})\bigotimes(x_{1},y_{1})$$M_{1}$$C_{1}$$(z_{2},M_{1})\bigotimes(x_{2},y_{2})$$M_{2}$$C_{2}$$M_{K}$$C_{K}$$(z_{K+1},M_{K})\bigotimes(x_{K+1},y_{k+1})$$M_{K+1}$$C_{K+1}$$\scriptstyle{Sz_{1}=P_{1}\oplus z_{1}}$$\scriptstyle{Sx_{1}=x_{1}*\delta_{1}^{2},Sy_{1}=y_{1}*\delta_{2}^{2}}$$\scriptstyle{Sz_{2}=P_{2}\oplus z_{2}}$$\scriptstyle{Sx_{2}=x_{2}*\delta_{1}^{2},Sy_{2}=y_{2}*\delta_{2}^{2}}\scriptstyle{......}$$\scriptstyle{P_{K+1}}$$\scriptstyle{Sx_{K+1}=x_{K+1}*\delta_{1}^{2},Sy_{K+1}=y_{K+1}*\delta_{2}^{2}}$
@@ 6.1.3 $Eagle^{*}$ decryption algorithm
For $Eagle^{*}$ decryptors, the ciphertext sequences are $C=($$C_{1}$, $C_{2}$, …, $C_{K+1}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1}$, $Sz_{1}$, $Sz_{2}$, …, $Sz_{K})$, $Su$ and $P_{K+1}$. Here, the ciphertext sequences can be divided into three independent parts $(C_{1},C_{2}$, …, $C_{K+1},Sz_{1},Sz_{2},...,Sz_{K},P_{K+1},Su)$ and $(Sx_{1},Sx_{2}$, …, $Sx_{K+1})$ and $(Sy_{1},Sy_{2}$, …, $Sy_{K+1})$, First we use $(\delta_{1})$ and $(Sx_{1},Sx_{2},...,Sx_{K+1})$ to calculate the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+1})$, then we use $(\delta_{2})$ and $(Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1})$ to calculate the random number sequences $y=(y_{1}$, $y_{2}$, …, $y_{K+1})$, then we use the random number $u=Su/\delta_{3}/\delta_{3}$ and the ciphertexts $(C_{1}$, $C_{2}$, …, $C_{K+1}$, $Sz_{1},Sz_{2},...,Sz_{K},P_{K+1})$ and the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+1})$ and $y=(y_{1}$, $y_{2}$, …, $y_{K+1})$ to decrypt $P=(P_{1}$, $P_{2}$, …, $P_{K})$. The specific process is detailed in the following algorithm.
@@
0:Key: $(\delta_{1},\delta_{2},\delta_{3})$. Ciphertext sequences:                 $(C_{1},C_{2},...,C_{K+1})$,                 $(Sx_{1},Sx_{2},...,Sx_{K+1})$,                 $(Sy_{1},Sy_{2},...,Sy_{K+1})$,                 $(Sz_{1},Sz_{2},...,Sz_{K})$,                 $Su,P_{K+1}$
@@
[…]
3:for $i=1;i\leq K+1;i++$ do
4:  $x_{i}\leftarrow(Sx_{i}/\delta_{1}/\delta_{1})\oplus 1$
5:  $y_{i}\leftarrow Sy_{i}/\delta_{2}/\delta_{2}$
6:end for
[…]
9:for $j=K;j\geq 1;j$-- do
10:  $C\leftarrow M/u/u$
11:  $C\leftarrow\psi(C)$
12:  $z_{j}\leftarrow\mathbb{D}_{x_{j},y_{j}}(C)$
13:  $M\leftarrow(C_{j}\oplus C)/(x_{j}\oplus y_{j})$
14:  $P_{j}\leftarrow z_{j}\oplus Sz_{j}$
15:end for
@@ 6.1.4 Basic Security Analysis
[Theorem 13] In the $Eagle^{*}$ encryption algorithm, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+1}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1}$, $Sz_{1}$, $Sz_{2}$, … , $Sz_{K}$, $Su$, $P_{K+1}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to cracking three respective keys $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+1})$ of the first A-Encrypt algorithm and only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+1})$ of the second A-Encrypt algorithm and only the ciphertext $Su$ of the third A-Encrypt algorithm.
@@
According to the definition of $Eagle^{*}$, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+1}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1}$, $Sz_{1}$, $Sz_{2}$, … , $Sz_{K}$, $Su$, $P_{K+1}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to solve the following system of equations.
@@
[…]
Where $x_{1},x_{2},...,x_{K+1},z,y_{1},y_{2},...,y_{K+1},u,\delta_{1},\delta_{2},\delta_{3}$ are unknown variables to be solved. $P_{1}$, $P_{2}$, …, $P_{K}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+1}$, $Sz_{1}$, $Sz_{2}$, …, $Sz_{K}$, $Su$, $P_{K+1}$ are known variables.
@@
The unknown variables $x_{1}$, $x_{2}$, …, $x_{K+1}$ in the system of equations (6.2) can be equivalently regarded as the plaintexts of the A-encryption algorithm, with the known variables $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+1}$ being the ciphertexts of the A-encryption algorithm, and $\delta_{1}$ being the key. Solving the system of equations (6.2) is equivalent to crack the key $\delta_{1}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+1})$ of A-Encrypt algorithm.
Similarly, solving the system of equations (6.3) is equivalent to crack the key $\delta_{2}$ under the condition of knowing only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+1})$ of the second A-Encrypt algorithm.
Similarly, solving the system of equations (6.4) is equivalent to crack the key $\delta_{3}$ under the condition of knowing only the ciphertext $Su$ of the third A-Encrypt algorithm. ∎
@@
(ii) Given any key, any plaintext in one group is equally possible.
(iii) In the case of knowing only ciphertext, any key for one group is equally possible.
According to the the Algorithm 11 and Algorithm 12, A-Encrypt obviously satisfies the above three conditions. This means that A-Encrypt can only use exhaustive search to crack the key when only knowing their ciphertexts, and there are no effective algorithms.
@@
It suffices to consider only the case where $\delta_{1}$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+1}$ can be directly calculated from the system of equations (6.2), but this does not change the conclusion that each equation in the system of equations (6.1) is an identity equation.
@@
When only exhaust $\delta_{3}$, since both $x_{i}$ and $y_{i}$ are unknown, it is impossible to determine any $M$ or $C$, which brings no benefit to solving the system of equations (6.1), (6.4).
@@
When enumerating over $(\delta_{1},\delta_{2})$, we can directly calcute $x_{1}$, $x_{2}$, …, $x_{K+1}$ from the system of equations (6.2), and directly calculate $y_{1}$, $y_{2}$, …, $y_{K+1}$ from the system of equations (6.3). Substituting these calculated $x_{i}$ and $y_{i}$ into the system of equations (6.1), we can sequentially calculate the hidden variables which related to $u$. Even we can solve $u$ quickly, we still need to solve the equation $u*\delta_{3}*\delta_{3}=Su$ to obtain $\delta_{3}$, this equation is a nonlinear equation about $\delta_{3}$. As long as Assertion 1 holds, the only way to solve this equation is exhaustively enumerating $\delta_{3}$.
@@
It suffices to consider only the case where $(\delta_{1},\delta_{3})$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+1}$ can be directly calculated from the system of equations (6.2). Since $y_{i}$ is unknown, it is impossible to determine any $M$ or $C$, which brings no benefit to solving the system of equations (6.1), (6.2), and (6.3).
@@
Due to the complete randomness of $x_{i}$ and $y_{i}$ and $z_{i}$, there is no statistical relationship or any form of mathematical relationship between the plaintext $P_{i}$ and $Sx_{i}$. Similarly, there is no statistical relationship between $P_{i}$ and $Sy_{i}$. Furthermore, according to Theorem 9, $P_{i}$ and $C_{i}$ are also completely unrelated. In fact, if we disregard all $M$ and $C$ (in reality, all $M$ and $C$ are invisible without solving for $x_{i}$ and $y_{i}$ and $u$), the encryption process of $Eagle^{*}$ is equivalent to one-time pad. An attacker cannot find any patterns between plaintexts and ciphertexts from a large sample of plaintext-ciphertext pairs.
After a simple deduction, we can obtain the following expression among plaintext, key, and ciphertext.
$P_{i+1}$
$=P_{i}*(Sx_{i}/\delta_{1}^{2}\oplus Sy_{i}/\delta_{2}^{2})$
$\oplus(Sx_{i}/\delta_{1}^{2})^{+(L-1)}*(Su/\delta_{3}^{2})$
$\oplus((Sx_{i+1}/\delta_{1})^{+(L-1)}\oplus C_{i+1})/(Sx_{i+1}/\delta_{1}^{2}\oplus Sy_{i+1}/\delta_{2}^{2})/(Su/\delta_{3}^{2})$
After collecting a large number of plaintext-ciphertext pairs, the attacker cannot directly decrypt new ciphertexts without cracking the key. This is because no matter what he does, he can only obtain the relationship between two adjacent pairs of plaintexts $P_{i+1}$ and $P_{i}$, and the parameters dependent on this relationship cannot be obtained solely based on the ciphertexts.